# Z-REx: Human-Interpretable GNN Explanations for Real Estate Recommendations

### v1

@@ Abstract.
Transparency and interpretability are crucial for enhancing customer confidence and user engagement, especially when dealing with black-box Machine Learning (ML)-based recommendation systems. Modern recommendation systems leverage Graph Neural Network (GNN) due to their ability to produce high-quality recommendations in terms of both relevance and diversity. Therefore, the explainability of GNN is especially important for Link Prediction (LP) tasks since recommending relevant items can be viewed as predicting links between users and items. GNN explainability has been a well-studied field, existing methods primarily focus on node or graph-level tasks, leaving a gap in LP explanation techniques.
This work introduces Z-REx, a GNN explanation framework designed explicitly for heterogeneous link prediction tasks. Z-REx utilizes structural and attribute perturbation to identify critical substructures and important features while reducing the search space by leveraging domain-specific knowledge. In our experimentation, we show the efficacy of Z-REx in generating contextually relevant and human-interpretable explanations for ZiGNN, a GNN-based recommendation engine, using a real-world real-estate dataset from Zillow Group, Inc. We also compare Z-REx to State-of-The-Art (SOTA) GNN explainers to show Z-REx’s superiority in producing high-quality human-interpretable explanations.
@@
copyright: noneconference: Knowledge Discovery and Data Mining; Due 10 February; TBDjournalyear: 2025ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Neural networks
@@ 2. Background and Related Works
GNN-based Recommendation Framework. Graph Neural Networks (GNNs) have emerged as a powerful approach for enhancing recommender systems. Over the past few years, several studies have applied GNNs directly to user-item bipartite graphs, yielding significant improvements in both effectiveness and efficiency (chen2020revisiting; he2020lightgcn; li2019hierarchical; sun2020neighbor; wang2019neural). One common challenge in this approach lies in capturing higher-order connectivity between nodes. Multi-GCCF (sun2019multi) and DGCF (liu2020deoscillated) address this by introducing artificial edges that connect two-hop neighbors (e.g., user-user and item-item graphs) to introduce proximity information into the user-item interaction. Node representations are computed layer-by-layer in GNNs, where the overall user and item representations are critical for downstream tasks like recommendation prediction. The most common practice is to adopt the final-layer embeddings as the ultimate representations (li2019hierarchical; ying2018graph). For a comprehensive survey of GNN-based recommender systems, readers are referred to (wu2022graph).
GNN-based recommender systems also differ in their node aggregation and update strategies. For example, NGCF (wang2019neural) enhances feature interactions between users and items using an element-wise product operation. NIA-GCN (sun2020neighbor), on the other hand, introduces pairwise neighborhood aggregation to better capture neighbor relationships. Inspired by GraphSAGE (hamilton2017inductive), (li2019hierarchical; sun2019multi; ying2018graph) utilize a concatenation operation followed by nonlinear transformations to update node representations. Conversely, LightGCN (he2020lightgcn) and LR-GCCF (chen2020revisiting) simplify the aggregation by removing non-linearities, which enhances both performance and computational efficiency.
@@ 4.2. Ground Truth
Negative examples are defined by selecting edges between users and cities (or listing) where no interaction has been recorded, assuming that non-interaction implies a lack of interest. These are weak negative edges, as there is no explicit evidence that the user dislikes the item. However, for our use case in real estate, we argue that a user’s lack of interaction and interest are closely aligned, given the context where the items are either listings or cities. Thus, incorporating weak negative edges to create a negative graph helps the recommendation model learn finer distinctions about the city (or listing) preferences and understand why users do not interact with them. A negative graph is generated by sampling negative edges from user-listing and user-city pairs as shown in Figure 2.
@@
We generate ground truth explanations by identifying the subgraphs most relevant for each prediction and compare their node features to identify their similarity. These subgraphs serve as interpretable justifications for why a specific user $u$ is likely to engage with listing $l$, thus providing a transparent and trustworthy recommendation system.
@@
Formally, let $\mathcal{Y}\subseteq\mathcal{V}_{u}\times\mathcal{V}_{l}$ be the set of observed interactions (positive examples), and $\mathcal{Y^{\prime}}\subseteq\mathcal{V}_{u}\times\mathcal{V}_{l}$ be the set of sampled negative examples. The training set $\mathcal{T}$ is constructed as $\mathcal{T}=\{(u,l,y_{ul})\mid(u,l)\in\mathcal{Y}\cup\mathcal{Y^{\prime}}$, where $y_{ul}=1$ if $(u,l)\in\mathcal{Y}$ and $y_{ul}=0$ if $(u,l)\in\mathcal{Y^{\prime}}$.
4.3. ZiGNN Architecture

### v2

@@ Abstract.
Transparency and interpretability are crucial for enhancing customer confidence and user engagement, especially when dealing with black-box Machine Learning (ML)-based recommendation systems. Modern recommendation systems leverage Graph Neural Network (GNN) due to their ability to produce high-quality recommendations in terms of both relevance and diversity. Therefore, the explainability of GNN is especially important for Link Prediction (LP) tasks since recommending relevant items can be viewed as predicting links between users and items. GNN explainability has been a well-studied field, but existing methods primarily focus on node or graph-level tasks, leaving a gap in LP explanation techniques.
This work introduces Z-REx, a GNN explanation framework designed explicitly for heterogeneous link prediction tasks. Z-REx utilizes structural and attribute perturbation to identify critical substructures and important features while reducing the search space by leveraging domain-specific knowledge. In our experimentation, we show the efficacy of Z-REx in generating contextually relevant and human-interpretable explanations for ZiGNN, a GNN-based recommendation engine, using a real-world real-estate dataset from Zillow Group, Inc. We compare against State-of-The-Art (SOTA) GNN explainers to show Z-REx outperforms them by 61% in Fidelity metric by producing superior human-interpretable explanations.
@@
copyright: rightsretainedconference: Knowledge Discovery and Data Mining Workshop on Machine Learning on Graphs in the Era of Generative Artificial Intelligence; August 06, 2025; Toronto, ON, Canadajournalyear: 2025ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Neural networks
@@ 1. Introduction
  • Z-REx outperforms PaGE-Link by 29%, GNNExplainer by 85%, and SubgraphX by 70% in the Fidelity metric, demonstrating its superior explanation accuracy.
@@ 2. Background and Related Works
GNN-based Recommendation Framework. Graph Neural Networks (GNNs) have emerged as a powerful approach for enhancing recommender systems. Over the past few years, several studies have applied GNNs directly to user-item bipartite graphs, yielding significant improvements in both effectiveness and efficiency (wang2023collaboration; mukherjee2023explaining; chen2020revisiting; sun2020neighbor; wang2019neural). One common challenge in this approach lies in capturing higher-order connectivity between nodes. Multi-GCCF (sun2019multi) and DGCF (liu2020deoscillated) address this by introducing artificial edges that connect two-hop neighbors (e.g., user-user and item-item graphs) to introduce proximity information into the user-item interaction.
Node representations are computed layer-by-layer in GNNs, where the overall user and item representations are critical for downstream tasks like recommendation prediction. The most common practice is to adopt the final-layer embeddings as the ultimate representations (li2019hierarchical; ying2018graph). For a comprehensive survey of GNN-based recommender systems, readers are referred to (wu2022graph).
NGCF (wang2019neural) enhances feature interactions between users and items using an element-wise product operation. NIA-GCN (sun2020neighbor), on the other hand, introduces pairwise neighborhood aggregation to better capture neighbor relationships. Inspired by GraphSAGE (hamilton2017inductive), (li2019hierarchical; sun2019multi; ying2018graph) utilize a concatenation operation followed by nonlinear transformations to update node representations. Conversely, LightGCN (he2020lightgcn) and LR-GCCF (chen2020revisiting) simplify the aggregation by removing non-linearities, which enhances both performance and computational efficiency.
@@ 4. Preliminaries
Src Node Type Dst Node Type Edge Type
User Listing views, saves, and tours
User City searched in
City Listing contains
Table 1. Description of relationships in the heterogeneous graph.
@@ 4.2. Ground Truth
Negative examples are defined by selecting edges between users and cities (or listing) where no interaction has been recorded, assuming that non-interaction implies a lack of interest. These are weak negative edges, as there is no explicit evidence that the user dislikes the item. However, for our use case in real-estate, we argue that a user’s lack of interaction and interest are closely aligned, given the context where the items are either listings or cities. Thus, incorporating weak negative edges to create a negative graph helps the recommendation model learn finer distinctions about the city (or listing) preferences and understand why users do not interact with them. A negative graph is generated by sampling negative edges from user-listing and user-city pairs as shown in Figure 2.
@@
We generate ground truth explanations by identifying the subgraphs most relevant for each prediction and comparing their node features to identify their similarity. These subgraphs serve as interpretable justifications for why a specific user $u$ is likely to engage with listing $l$, thus providing a transparent recommendation system.
@@
Let, $\mathcal{Y}\subseteq\mathcal{V}_{u}\times\mathcal{V}_{l}$ be the set of observed interactions (positive examples), and $\mathcal{Y^{\prime}}\subseteq\mathcal{V}_{u}\times\mathcal{V}_{l}$ be set of sampled negative examples. The training set $\mathcal{T}$ is constructed as $\mathcal{T}=\{(u,l,y_{ul})\mid(u,l)\in\mathcal{Y}\cup\mathcal{Y^{\prime}}$, where $y_{ul}=1$ if $(u,l)\in\mathcal{Y}$ and $y_{ul}=0$ if $(u,l)\in\mathcal{Y^{\prime}}$.
Figure 3. Z-REx overview.
5. Z-REx Overview
The workflow of Z-REx is illustrated in Figure 3 and algorithm is described in Algorithm 2 where Z-REx first performs feature perturbation to find the important features and then performs graph structural perturbations to identify the important subgraph structures. We focus on user-city due to the problem definition in § 3, but this method works for any relationship.
5.1. Feature Perturbation.
To interpret the GNN-based recommendation system, we first analyze the influence of target node features using feature perturbation. Consider a user $u$ and their recommended city $c_{t}$ with their embeddings, $h_{u}$ and $h_{c_{t}}$, where cosine similarity represents the predicted affinity score that the GNN uses to rank city $c_{t}$ for user $u$:
The feature perturbation involves the following steps:
  1. (1)
  2. (2)
5.2. Structural Perturbation.
While feature perturbation focuses on node-level characteristics, structural perturbation analyzes the impact of graph topology on the model’s predictions. Thus, accounting for the whole graph context in each recommendation explanation. Please note that structural perturbation happens with only the subset of features identified in the previous step. Feature perturbation identifies node attributes that have a significant influence on the recommendation outcomes, while structural perturbation uncovers graph edges and relationships that are critical to the model’s predictions.
  1. (1)
  2. (2)
    Identify Co-clicked Cities: For the subgraph $\mathcal{G}_{u}^{k}$, we identify pairs of cities $(c_{i},c_{j})$ that share a common predecessor user $u_{p}$. These pairs are added as new edges, representing co-click relationships, and their contributions to the model predictions are evaluated. By co-clicked cities, we mean cities in which the current user clicked listings, as well as cities other users clicked listings in, where they share at least one common city with the current user. A co-clicked city indicates user groups with similar preferences.
  3. (3)
Edges with the highest absolute $\Delta\text{sim}$ values are identified as critical contributors to the recommendation. These edges represent strong graph relationships that drive user preferences for specific cities. The hyperparameters that influence the structural perturbations are: (1) $k$ (hop distance) as it limits the number of edges perturbed to efficiently evaluate while focusing on the most impactful connections and increasing $k$ captures more indirect relationships but may introduce noise, and (2) edge removal strategy as prioritizing edges based on shared predecessors ensures that only influential connections are analyzed.
5.3. Graph Size vs. Resource Overhead
We provide real estate-specific recommendation explanations by leveraging the insight that co-clicked cities have a significant impact on recommendations. Co-clicked cities restrict the search space for structural perturbations, which in turn lead to finding the optimal number of edges required to change the recommendations. We acknowledge that iterative perturbation can be a resource-intensive process. But, our experiments show that real estate interaction is sparse in nature (as seen in Table 7), since people view the most number of listings, but only save a few, and tour even fewer listings. Therefore, the interaction graph from the real-world data is also sparse (as seen in the 30-days sub-section of Table 7), so large-scale real-world deployments are feasible.
As seen in Table 8 in evalaution, training one epoch on a 30-day interaction graph containing 29.8M edges requires 10.89 sec, only an order of magnitude slower than the 0.88 sec required for a 3-day snapshot containing 1.2M edges, confirming that additional temporal coverage does not explode computation and inference time stays almost the same. The results demonstrate that Z-REx delivers timely and faithful explanations at a production scale.
5.4. Generalizability of Z-REx
To adapt the Z-REx for other recommendation tasks, the edge perturbation step needs modification since for real estate, we are using domain knowledge and are focusing on co-clicked edges. For other domains, these edges might not exist, or the assumption we made might not be accurate. Since the structure influences the GNN’s message passing step, we think that for other domains, we can translate the insights to structural constraints, we can update § 5.2 Structural Perturbation, and decrease the search space from all the potential edges to a smaller list of potentially important edges.
6. Evaluation
6.1. Methodology
The data processing of ZiGNN is described in Algorithm 3, which begins by extracting features containing user activities (e.g., views and saves), user and listing attributes, and their geographical regions. During preprocessing, missing values are handled systematically to ensure data integrity. Outliers, which can distort analysis and model training, are addressed by replacing extreme values (detected using Z-scores) with the column mean. These steps ensure a clean and reliable dataset. Additionally, numerical features are normalized using z-score normalization, creating a consistent data representation. These preprocessing steps create a robust and standardized dataset, ready for graph-based modeling in the recommendation system.
After preprocessing, the user, listing, and city data are used to construct a heterogeneous graph. Each entity (e.g., user, listing, and city) is represented as a unique node type, and edges capture interactions between these nodes. For instance, user-to-listing edges represent activities such as views or saves, while city-to-listing edges denote listings contained within a city. Node features are assigned and this comprehensive graph representation forms the backbone of the recommendation system, enabling the ZiGNN to understand complex relationships.
The training pipeline first constructs the negative graph to ensure a balanced training dataset by mimicking real-world scenarios where users do not interact with all items. During evaluation, the ZiGNN’s effectiveness was tested using node embeddings derived from the trained model. These embeddings were normalized to enable cosine similarity-based retrieval, crucial for recommendation tasks. For specific canonical edge types (e.g., ’user’, views, ’city’), embeddings were computed separately for source (e.g., user) and destination nodes (e.g., city). The recommendation performance of ZiGNN is measured using nDCG@K, which evaluates the ranking quality of recommendations and prioritizes highly relevant items. Metrics like nDCG@K was used to evaluate recommendation quality, providing quantitative insights.
Metric. Z-REx explains the recommendations generated by ZiGNN, and for this experimentation, we focus on user-to-city views relationships, but this works for any relationships. First, we identify an important subset of features by measuring the change in nDCG@K score when individual features are zeroed out. Then, selecting those that produce a negative impact on nDCG, meaning those features were important for predicting relevant recommendations. To align with the recommendation explainability task, we redefine the traditional Fidelity metric (yuan2021explainability)—originally based on change in prediction confidence—as the change in nDCG@K score due to perturbations. The explanation performance of Z-REx is then evaluated based on the drop in nDCG@K score and the change in cosine similarity ($\Delta\text{sim}$) between user and city embeddings caused by structural perturbations and by only using the important features.
6.2. Dataset
The dataset was collected over a three-day and 30-day period from the state of Washington in the USA. The dataset statistics are shown in Table 5 and Table 6, and more details in § A.4. The dataset is split into training, validation, and test sets, ensuring that the test set contains previously unseen pairs, which are evaluated on the following day to avoid any data leakage. The 30-day setting naturally reduces temporal leakage between training and evaluation interactions, further strengthening the case for Z-REx as a robust, scalable, and domain-aware explainer for real-estate recommender systems. Its ability to combine high fidelity with diversified recommendations makes it the only method that meets practical requirements for accuracy, interpretability, and user engagement.
To note, currently there are no public real estate datasets that contain user-to-listing interactions or the rich attributes we are using. Therefore, we collected and used our own dataset, that accounts for seasonal trends. In the 3-days training data (refer to Table 7), expected trends are seen such as view events are the most popular event, followed by save and then toured because users tend to view multiple items before saving them and finally touring the item. The average interaction of the user to listing for view event for the 25th quantile is 1.0 and the 75th quantile is 8.0, the save event for the 25th quantile is 2.7 and the 75th quantile is 3.0, and the tour event for the 25th quantile is 3.2 and the 75th quantile is 4.0. Therefore, from the average interaction of different events statistics tells us that the graph generated from the dataset will contain more sparse connections than dense connections.
Figure 4. Performance of ZiGNN against different baselines.
6.3. Recommendation Accuracy of ZiGNN
To investigate RQ1, we compared the performance between the ZiGNN (3-days) and the baseline recommendation models to reveal ZiGNN consistently and substantially outperforms the rest as seen in Figure 4. ZiGNN consistently surpasses the histogram model across all values of K, with the performance gap widening as K increases. Notably, at $K=10$, the ZiGNN achieves an nDCG score of 0.019, a 52% improvement over the histogram model’s 0.0125, demonstrating ZiGNN’s ability to produce more accurate and relevant recommendations as the recommendation set grows.
This trend is further emphasized by the steep growth curve of the GNN’s performance compared to the relatively linear increase observed for the histogram model, showcasing the ZiGNN’s adaptability. ZiGNN’s structural representation learning captures nuanced relationships in the data that the histogram model, with its simpler statistical approach, fails to address. At $K=1$, where only the top recommendation matters, ZiGNN outperforms the histogram model by a remarkable 200% (0.0075 vs. 0.0025), highlighting its capacity to prioritize the most relevant results effectively. Comparing ZiGNN performance using the extended 30-day dataset and 3-day dataset, the 30-day model yields consistent improvement gains ranging from  30% at $K=10$ to nearly 90% at $K=1$, highlighting the benefit of incorporating a longer 30-day interaction window where the user behavior is captured with higher relevance.
Figure 5. Impact of zeroing out features to find important features.
Figure 6. Impact of sequentially adding important features on nDCG.
6.4. Explanation Quality of Z-REx
For investigating RQ2, we first consider the feature perturbation quality and show that we can identify the top ten important city features for the recommendation by measuring the difference between the original nDCG and the perturbed nDCG when a feature is zeroed out. When the difference is positive, the feature was important for predicting the recommendation, as shown in Figure 5, and features with negative differences mean they were confusing the model (more details in § A.5).
In particular, these important features can be grouped: (1) geographic - latitude (top right and bottom left) and longitude (bottom left), (2) boolean - view, spa, and vacant land, and (3) numeric - year, population count, days on the market, and square feet. As illustrated in Figure 6, using only the important features yield better performance than using all the features, demonstrating that Z-REx effectively identifies the important features needed for user-city recommendation. These top features will be the only features used for measuring the structural perturbation quality.
The important city features align with domain knowledge: a city is typically characterized by its geographic location and aggregate properties (e.g., population count, the average year of houses, and average days on the market), which are different from the features usually used to characterize a listing (e.g., the number of bathrooms and bedrooms). Aggregate properties offer a holistic snapshot of a city’s dynamics and freshness. Interestingly, cities with a younger infrastructure experience rapid growth, indicating a strong interplay between user preference and given recommendations.
3-days (Training)
nDCG (decr.) 94% 81% (-13%) 21% (-73%) 47% (-47%)
cosine similarity -0.10 -0.07 (-0.03) -0.02 (-0.08) -0.04 (-0.06)
30-days (Training)
nDCG (decr.) 92% 63% (-29%) 9% (-85%) 22% (-70%)
cosine similarity -0.09 -0.05 (-0.04) -0.01 (-0.08) -0.02 (-0.07)
Table 2. Quantitative evaluation of Z-REx against GNN explainers.
Our month-long (30-days) evaluation confirms that Z-REx delivers markedly stronger explanations than all baselines while optimally scaling to a larger interaction graph. With a 30-days training window, Z-REx preserves 92% of the original nDCG, almost identical to its 3-day performance (94%), and maintains the largest embedding shift (-0.09), indicating that it highlights influential edges without collapsing recommendation diversity.
The stability of Z-REx across time windows suggests that its structural perturbation strategy generalizes rather than over-fitting to short-term interaction noise. The two-point drop in nDCG retention and the 0.01 improvement in cosine similarity from 3 to 30 days imply that extending the past horizon yields the same explanatory fidelity with slightly more conservative embedding movements, an attractive trade-off for production systems.
6.5. SOTA GNN Explainers vs. Z-REx
We compared Z-REx against a specialized GNN-based recommendation explainer PaGE-Link (zhang2023page) and two general GNN explainers, GNNExplainer (ying2019gnnexplainer) and SubgraphX (yuan2021explainability) in Table 2 to answer RQ3. Z-REx performed the best, followed by PaGE-Link and general purpose GNN explainers performed the worst which is expected since Z-REx and PaGE-Link have been specifically designed for LP tasks. In contrast, the recommendation-specific PaGE-Link retains only 63% of nDCG and induces a much smaller change in cosine similarity, while general-purpose explainers fall to 22% (SubgraphX) and 7% (GNNExplainer). Competing methods degrade sharply because their mask-optimization (GNNExplainer) or Monte-Carlo subgraph search (SubgraphX) faces an exponentially expanding candidate space as graph size grows.
Since, PaGE-Link creates the k-hop ego-graph by limiting the consideration of relevant features and subgraphs. Therefore, PaGE-Link performance degrades sharply because they compute the k-hop neighborhood around the target link, but there is no guarantee that the k-hop relationship will include relevant links. This limitation is reflected by -0.04 less decrease is cosine similarity. Z-REx avoids this pitfall by exploiting domain knowledge: the number of co-clicked cities grows sub-linearly, so the search space for meaningful perturbations remains tractable even on month-long graphs. Compared to general purpose GNN explainer, i.e., GNNExplainer and SubgraphX, Z-REx demonstrates average of -78% more reduction in nDCG and -0.08 more cosine similarity increase, indicating that it more effectively identifies the features and subgraphs the GNN relies on to make recommendations.
Hyperparameter Values Tested nDCG@1 Scores Best Score
Neighborhood Size ($k$) 4, 8, 16 0.0070, 0.0075, 0.0071 0.0075 (at $k{=}8$)
Edge Removal Strategy PRI, HID, HBC 0.0070, 0.0077, 0.0075 0.0077 (at HID)
Negative Graph Size 1, 3, 5 0.0072, 0.0074, 0.0075 0.0075 (at 5)
Learning Rate $10^{-3}$, $10^{-2}$, $10^{-1}$ 0.0072, 0.0074, 0.0069 0.0074 (at $10^{-2}$)
Output Feature Dim. 128, 256, 512 0.0071, 0.0075, 0.0081 0.0081 (at 512)
Table 3. Hyperparameter sensitivity summary. Each row varies one hyperparameter while holding the others constant; the rightmost column reports the highest nDCG@1 observed for that setting.
6.6. Hyperparameter Sensitivity Study
Table 3 shows the hyperparameter sensitivity tuning summary where we test different neighborhood sizes ($K$), edge removal strategy, size of negative graph, learning rate and output feature dimensions. The size of the input feature dimension is governed by the number of features for each entity type, so we did not incorporate that. The three different edge removal strategies we used are: (1) PRI (prioritize Recent Interaction) which removes edges with more recent interactions first (reduces the strength of fresh user-item relationships); (2) HID (High In-Degree) which removes edges to nodes with the highest in-degree (breaking these highly connected links disrupts potential shared neighborhoods); and (3) HBC (High Betweenness Centrality) which removes edges with high betweenness centrality, targeting the links acting as bridges in the hub cities in the interaction graphs.
Increasing the neighborhood size ($K$) from 4 to 8 improves nDCG@1, but going further to 16 diminishes it, indicating that while broader neighborhoods capture more context, too large a neighborhood may dilute relevant signals. HID (high in-degree) as the edge removal strategy yields the best score among the tested strategies, suggesting that selectively removing overly connected edges can help avoid excessive overlap and improve the model’s focus on more distinct interactions. A larger negative graph size, up to a point, seems beneficial (optimal at size 5 in this table) because it gives the model a more balanced signal to distinguish relevant versus irrelevant connections. Tuning the learning rate reveals that a rate of 1e-2 provides a stable convergence path. Finally, a higher output feature dimension (512) significantly boosts nDCG@1, demonstrating that a richer embedding space helps the model capture more nuanced relational patterns but at the cost of resource consumption.
Figure 7. Z-REx’s explanation of a recommended city #1.
7. Case Study
In the case study, we selected a real user (i.e., customer) and, based on their interactions, generated a city recommendation along with explanations. To enhance visualization, we present a simplified graph (Figure 7) focusing on the customer (yellow node), co-clicked cities (blue nodes), and the recommended city (green node). A blue edge will exist between the co-clicked city nodes, the recommended city is the green node, and the red edges are the edges identified as important for recommendations.
8. Conclusion
We introduced Z-REx, which aims to bridge a critical gap in explainable ML for recommendation systems by providing human-interpretable insights into GNN-based recommendation systems to increase user confidence and customer engagement. Z-REx offers contextually relevant and human-interpretable explanations through structural and attribute perturbation. Z-REx outperforms PaGE-Link by 29%, GNNExplainer by 85%, and SubgraphX by 70% in the Fidelity metric, demonstrating its superior explanation accuracy compared with the ground-truth. Specifically, ZiGNN provides whole graph instance-level explanations for GNN models in the context of the heterogeneous link prediction (LP) task by identifying a subset of node features and subgraphs that are most influential for GNN’s-based recommendation system.
[…]
[heading=none]
Appendix A Appendix
A.1. ZiGNN Architecture